# CGKN: A Deep Learning Framework for Modeling Complex Dynamical Systems and Efficient Data Assimilation

### v1

@@ 4 Discussion and Conclusion
In this work, we propose a conditional Gaussian Koopman network — CGKN — as a unique deep learning framework for modeling complex dynamical systems that facilitate both accurate forecasts and efficient DA. The proposed framework is inspired by the Koopman theory and builds upon the conditional Gaussian nonlinear system, transforming any nonlinear system into a modeled system that embeds a conditional linear structure to facilitate efficient analytic formulae of DA. The proper nonlinear transformation is jointly learned with the unknown terms in the transformed conditional Gaussian nonlinear system, based on a total loss function that accounts for the overall performance of the nonlinear transformation, the state prediction, and the DA. Among all the models that support efficient analytic formulae of DA, the numerical examples demonstrate that the CGKN model generally outperforms the simplified KoopNet model and the CG-Reg model in both state prediction and DA.
The proposed framework integrates the strengths of deep learning methods, Koopman theory, and conditional Gaussian nonlinear systems for modeling complex dynamical systems and efficient DA. In addition, it serves as an effort towards a systematic approach to impose useful structures into deep learning models and to facilitate their usage in other outer-loop applications of scientific machine learning, e.g., control of partially observable systems, and inverse problems. Some extensions and future directions of the proposed framework include more rigorous approaches to efficiently quantify the uncertainties associated with the DA results of the CGKN model and multi-objective optimization methods to enhance the overall performance of the nonlinear transformation, the state predictions, and the DA.

### v2

@@ 4 Discussion and Conclusion
In this work, we propose a conditional Gaussian Koopman network — CGKN — as a unique deep learning framework for modeling complex dynamical systems that facilitate both accurate forecasts and efficient DA. The proposed framework is inspired by the Koopman theory. It builds upon the conditional Gaussian nonlinear system, transforming any nonlinear system into a modeled system that embeds a conditional linear structure to facilitate efficient analytic formulae of DA. The proper nonlinear transformation is jointly learned with the unknown terms in the transformed conditional Gaussian nonlinear system, based on a total loss function that accounts for the overall performance of the nonlinear transformation, the state prediction, and the DA. Among all the models that support efficient analytic formulae of DA, the numerical examples demonstrate that the CGKN model generally outperforms the standard KoopNet model and the CG-Reg model in both state prediction and DA.
The proposed framework integrates the strengths of deep learning methods, Koopman theory, and conditional Gaussian nonlinear systems for modeling complex dynamical systems and efficient DA. In addition, it serves as an effort towards a systematic approach to impose useful structures into deep learning models and to facilitate their usage in other outer-loop applications of scientific machine learning, e.g., control of partially observed systems, and inverse problems. Some extensions and future directions of the proposed framework include more rigorous approaches to efficiently quantify the uncertainties associated with the DA results of the CGKN model and multi-objective optimization methods to enhance the overall performance of the nonlinear transformation, the state predictions, and the DA.
@@ Acknowledgments
Appendix A Additional Results of the Projected Stochastic Burgers–Sivashinsky Equation
Table A.1 summarizes some additional results of the projected stochastic Burgers-Sivashinsky equation, for comparing the performance of CGKN and other methods with different choices of observed and unobserved state variables. In addition to the setup of state variable $x$ being observed as studied in Section 3, we also investigate two different setups with state variables $y$ and $z$ being observed, respectively. The results in Table A.1 confirm that the performance of CGKN is not sensitive to the specific choice of observed and unobserved states presented in Section 3.
Table A.1: Test results of state forecast and data assimilation under different setups of observed and unobserved states for the example of projected stochastic Burgers–Sivashinsky equation. The errors are the normalized root mean squared error (NRMSE) between true values and approximated values.
Models Errors Obs:$x$     Unobs:$y,z$ Obs:$y$     Unobs:$x,z$ Obs:$z$     Unobs:$x,y$ Forecast Error DA Error Forecast Error DA Error Forecast Error DA Error True Model 2.4566e-01 6.9466e-01 2.4566e-01 9.7836e-01 2.4566e-01 1.4051e+00 CGKN 2.8389e-01 7.2776e-01 2.9901e-01 1.0992e+00 2.9863e-01 8.4302e-01 Standard KoopNet 3.7600e-01 7.5069e-01 4.1460e-01 1.3753e+00 3.3586e-01 8.8313e-01 CG-Reg 2.7766e-01 1.2884e+00 3.3051e-01 1.3919e+00 4.4084e-01 6.5044e+00 DNN 2.4728e-01 2.4728e-01 2.4728e-01
Appendix B A Quick Summary of the ENSO SPDE Test Model
The physics-based reference ENSO model [113] discussed in this paper is built based on the dynamics of atmosphere-ocean interactions in the tropical Pacific. The model consists of two main components: a deterministic dynamical core that captures the essential physics of ENSO and stochastic parameterizations that trigger the extreme events showing the atmospheric and oceanic processes occurring at different timescales.
The dynamical core describes the interactions between the atmosphere, the ocean, and the SST through three key mechanisms. First, the atmosphere drives ocean circulation through wind stress, where the wind busts generate currents and affect the thermocline depth. Second, the ocean affects the atmosphere through latent heat, which is proportional to SST. Third, the ocean modulates SST through changes in the thermocline depth and zonal currents.
The atmospheric component is based on a non-dissipative Matsuno-Gill type atmosphere model  [116, 117]. This model is a simplified representation of atmospheric dynamics in the tropics, designed to capture large-scale motions. A key feature in the model of Eq. (B.1) is the latent heating $E_{q}$, which is proportional to the sea surface temperature and drives the atmospheric circulation. For the ocean component, the model employs a simple shallow-water system  [118]. The model focuses on the dynamics of an active upper layer. The upper layer dynamics are driven by surface wind stress from the atmosphere. This wind stress causes currents and waves, representing momentum transfer from the atmosphere to the ocean. These assumptions result in the reduced shallow-water equations shown in Eq. (B.2). The SST equation is derived from an SST budget equation  [119]. These include vertical temperature changes due to thermocline variations, horizontal heat transport by ocean currents, and thermal damping through air-sea heat exchange. These essential mechanisms are captured in Eq. (B.3).
The dynamical core is a coupled atmosphere-ocean-SST system:
Atmosphere:
$-yv-\partial_{x}\theta=0$ (B.1)
$yu-\partial_{y}\theta=0$
$-\left(\partial_{x}u+\partial_{y}v\right)=E_{q}/(1-\bar{Q})$
Ocean:
$\partial_{t}U-c_{1}YV+c_{1}\partial_{x}H=c_{1}\tau_{x}$ (B.2)
$YU+\partial_{Y}H=0$
$\partial_{t}H+c_{1}\left(\partial_{x}U+\partial_{Y}V\right)=0$
SST:
$\partial_{t}T=-c_{1}\zeta E_{q}+c_{1}\eta_{1}H+c_{1}I\eta_{2}U.$ (B.3)
In the atmosphere component (B.1), $u$, $v$, and $\theta$ represent the zonal wind speeds, meridional wind speeds, and the potential temperature, respectively. In the ocean components (B.2), $U$, $V$ are zonal and meridional ocean currents, $H$ is the thermocline depth. Besides, $T$ represents the SST. These variables are all anomalies. The $t$ represents the interannual time coordinate, $x$ is the zonal coordinate, and $y$ and $Y$ are the meridional components for the atmosphere and ocean components. The equations include various parameters such as latent heat $E_{q}=\alpha_{q}T$, wind stress $\tau_{x}$, latent heating exchange coefficient $\zeta$, and the strengths of the thermocline $\eta_{1}$ and zonal advective feedback $\eta_{2}$. The relative magnitudes of $\eta_{1}$ and $\eta_{2}$ differ between the Eastern Pacific (EP) and Central Pacific (CP) due to variations in thermocline depth and zonal SST gradients. The model assumes periodic boundary conditions for the atmosphere and reflection boundary conditions for the Pacific Ocean. Since ENSO is primarily a phenomenon near the equator, we further project the model to the leading basis in the meridional direction.
While this deterministic dynamical core captures the basic ENSO physics, additional processes at different timescales play crucial roles in generating realistic features of ENSO diversity. These additional effects can be effectively characterized by stochastic parameterizations, which describe the intraseasonal wind variations and the decadal changes in the background conditions.
The wind stress $\tau_{x}$ consists of two components, the atmospheric circulation $u$ and the stochastic wind bursts $u_{p}$, which is localized in the western Pacific (WP) and has the following structure:
$u_{p}(x,y,t)=a_{p}(t)s_{p}(x)\phi_{0}(y)$ (B.4)
where $\phi_{0}(y)$ is the leading meridional basis, $s_{p}(x)$ is a fixed spatial structure, and $a_{p}(t)$ is the wind burst amplitude governed by a stochastic process:
$\frac{\mathrm{d}a_{p}}{\mathrm{~{}d}t}=-d_{p}a_{p}+\sigma_{p}(T_{C})\dot{W}_{p}$ (B.5)
Here, $d_{p}$ is the damping term, $\dot{W}_{p}$ is a white noise source, and $\sigma_{p}(T_{C})$ is the state-dependent noise strength, which is a function of the SST averaged over the CP. In the absence of seasonal cycle and decadal influence, the noise strength is parameterized as $\sigma_{p}(T_{C})=1.6(\tanh(T_{C})+1)$. This state-dependent noise coefficient implies that wind burst activity is more active during El Niño events due to the eastward extension of the warm pool. The intraseasonal wind bursts not only affect the interannual variability but are also modulated by it.
Finally, the decadal variability is driven by a simple stochastic process:
$\frac{\mathrm{d}I}{\mathrm{~{}d}t}=-\lambda(I-m)+\sigma_{I}(I)\dot{W}_{I},$ (B.6)
where $\lambda$ is the damping term set to 5 years-1, representing the decadal time scale. $\sigma_{I}(I)$ is the state-dependent noise strength, and $\dot{W}_{I}$ is the white noise source. The state-dependent noise coefficient allows the distribution of $I$ to be non-Gaussian, with $I$ representing the strength of the easterly trade wind in the lower level of the Walker circulation on the decadal time scale. A uniform distribution between $[0,1]$ is adopted for $I$, with larger values corresponding to stronger easterly trade winds.